OpenStudy (anonymous):

PLEASE HELP!!!!!! For the functions f(x) = 2x − 6 and g(x) = 5x + 1, which composition produces the greatest output?

ganeshie8 (ganeshie8):

got options ?

OpenStudy (anonymous):

A. Both compositions produce the same output. B. Neither composition produces an output. C. f(g(x)) produces the greatest output. D. g(f(x)) produces the greatest output.

OpenStudy (anonymous):

@Here_to_Help15 @UsukiDoll

rvc (rvc):

@Michele_Laino please help us :)

ganeshie8 (ganeshie8):

Maybe lets work f(g(x)) and g(f(x))

ganeshie8 (ganeshie8):

f(x) = 2x − 6 and g(x) = 5x + 1 f(g(x)) = ?

OpenStudy (anonymous):

10x - 4?

ganeshie8 (ganeshie8):

g(f(x)) = ?

OpenStudy (anonymous):

10x - 29

ganeshie8 (ganeshie8):

Excellent! so we have f(g(x)) = 10x-4 g(f(x)) = 10x-29

ganeshie8 (ganeshie8):

Notice that g(f(x)) = 10x - 29 = 10x - 4 - 25 = f(g(x)) - 25

ganeshie8 (ganeshie8):

which is same as f(g(x)) = g(f(x)) + 25

ganeshie8 (ganeshie8):

can we say f(g(x)) is always 25 greater than g(f(x)) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it would be C?

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

thank you :)

ganeshie8 (ganeshie8):

yw